OpenStudy (anonymous):

How do I simplify 3a/a+3 - (a-4)/a-2

OpenStudy (anonymous):

Send the (a-4)/(a-2) to the other side. 3a/a+3 = (a-4)/(a-2) 3a/a + 3a/a = (a-4)/(a-2) (3a+3)/ a = (a-4)/(a-2) Then cross multiply. (3a+3)(a-2) = (a-4)(a) Solve. 3a^2 - 6a - 6 = a^2 -4a Make everything come to one side. 2a^2 - 2 - 6 = 0 Factorise. (2a +2)(a-3)= 0 Thus Either a = -1 or a= 3